OpenStudy (anonymous):

Show all work to simplify 1 over x plus 3 over 8 = 1 over 2. Use complete sentences to explain how to simplify this expression.

OpenStudy (harsimran_hs4):

@BelleWalker we have a dedicated equation generator button below...please use it for better understanding of question to others and avoid any confusion

OpenStudy (anonymous):

\[\frac{ 1 }{ x } + \frac{ 3 }{ 8 } = \frac{ 1 }{ 2 } \] I am sorry about that!

OpenStudy (harsimran_hs4):

lets begin 1. take the lcm and obtain a common denominator \[\frac{ 8 + 3x}{ 8x } = \frac{ 1 }{ 2 }\]

OpenStudy (harsimran_hs4):

2. rearranging... 16 + 6x = 8x now can you solve for x?

OpenStudy (anonymous):

Yes thank you so much!

OpenStudy (harsimran_hs4):

:)